Circular No. 6812

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998C IN UGC 3825 Q. Y. Qiao, Y. L. Qiu, W. D. Li and J. Y. Hu, Beijing Astronomical

Observatory (BAO) report their discovery of a supernova in UGC 3825 on unfiltered CCD images taken with 0.6-m telescope on Jan. 21 UT. SN 1998C is located at α = 7^{h}23^{m}34^{s}.63, δ = +41^{o}26′04′′.2 (equinox 2000.0), which is 15′′.5 east and 0′′.5 south of the center of UGC 3825. CCD magnitudes: Jan. 21.57, 18.1; 22.59, 17.9. Unfiltered CCD images of the same field taken on Jan. 8 show no star at this position (limiting mag ∼ 18.5). A lowdispersion spectrum of the supernova taken on Jan. 22.70 by Qiao and X. Y. Dong (BAO) and X. Zhang (Yunnan Astronomical Observatory) with the BAO 2.16-m telescope at Xinglong Station shows this to be a type-II supernova before maximum. The spectrum displays a blue continuum with broad emission lines of Hα and Hβ superimposed. The FWHM of Hα shows an expansion velocity of 8500 km/s. The recession velocity of UGC 3825 derived from the narrow Hα emission line is 8600 km/s.

COMET C/1995 O1 (HALE-BOPP) G. Garradd, Loomberah, N.S.W., reports that his CCD images taken

from the second half of 1997 Dec. through 1998 Jan. 17.46 UT show an antitail, and the December images additionally show a longer normal dust tail than previously; a Jan. 4 image can be viewed at http://usrwww.mpx .com.au/~gjg/halebop.htm. On Jan. 6.70, he found the antitail 1^{o}.5 long visually with a 0.25-m reflector, much fainter than the main dust tail (which was seen to length 4^{o}.5 with 10×50 binoculars).

H. Boehnhardt, O. Hainaut, G. Pizarro, and R. West, European Southern Observatory, write: “Schmidt telescope exposures of comet C/1995 O1 taken between 1997 Dec. 29 and 1998 Jan. 6 at La Silla show a straight tail ∼ 10′′ wide that extends both 25′ into the sunward side of the nucleus at p.a. 8^{o} and at least 4^{o} to the anti-solar side at p.a. 188^{o}. This needle-like tail feature partially overlaps with the normal diffuse dust tail northward of the nucleus. The tail/antitail phenomenon is interpreted as due to the presence of large and old dust grains in the cometary dust tail that were released by the nucleus at least 100 days before the observations (or even much earlier).”

Visual m_{1} estimates (B = binoculars): Jan. 6.76 UT, 7.8 (A. Pearce, Nedlands, W. Australia, 20×80 B); 17.06, 8.0 (R. Lourençon, Jundiai, Brazil, 20×80 B); 21.4, 7.7 (S. O’Meara, Volcano, HI, 10×50 B).

1998 January 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green